i am helping my son with a worksheet from school and the answer sheet for his worksheet is saying his answer is wrong the question he is stuck on is (2xy)(3xy)(6x^8y^6) he is doing a multiplication of monomials worksheet and the answer he got is 36x^10y^6 is that right or is the worksheet right when it states the answer is 11x^10y^6 and why?
Answer by jim_thompson5910(28536)

The answer is actually 36x^10y^8, so the worksheet is wrong.


The final coefficient is equal to the product of the other coefficients. So the final coefficient is 2*3*6=6*6=36


Then you multiply the x terms to get x*x*x^8 = x^(1+1+8)=x^10

and then you multiply the y terms to get y*y*y^6 = y^(1+1+6) = y^8


put this all together to get 36x^10y^8